OpenStudy (anonymous):

Namelss' Riddle: Arrange the numbers 1 3 4 6 in an equation to make 24. Each number must be used once and only once. You can use any number and combination of addition, subtraction, multiplication or division operators, and brackets to enforce an order of evaluation. e.g. 4 + ((3 - 1) x 6), wrong result, but valid form. No tricks: - It's 24 in decimal - Each number is separate, you can't use 1 and 4 together to make fourteen etc. - It's not a word game where 24 means "to for" or "a day" or anything like that.

OpenStudy (anonymous):

because they were grandfather-father-son.

OpenStudy (anonymous):

Wow it took me like and hour to figure this out at first. Good job. Answer(as mentioned):one of the 'fathers' is also a grandfather. Therefore the other father is both a son and a father to the grandson. In other words, the one father is both a son and a father.

OpenStudy (anonymous):

Thank you for the medal<3

OpenStudy (anonymous):

Arrange the numbers 1 3 4 6 in an equation to make 24. Each number must be used once and only once. You can use any number and combination of addition, subtraction, multiplication or division operators, and brackets to enforce an order of evaluation. e.g. 4 + ((3 - 1) x 6), wrong result, but valid form. No tricks: - It's 24 in decimal - Each number is separate, you can't use 1 and 4 together to make fourteen etc. - It's not a word game where 24 means "to for" or "a day" or anything like that.

OpenStudy (anonymous):

Are exponents allowed?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

only standard operators.. no algebra either :)

OpenStudy (andriod09):

Can you an operation more than once?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

well yes you would hav to

OpenStudy (anonymous):

sorry misread

OpenStudy (anonymous):

yall want me to just post the answer? i gotta leave soon :)

OpenStudy (andriod09):

no the answer is

OpenStudy (andriod09):

sure i dont got the answer

OpenStudy (anonymous):

I'm pretty stumped too.

OpenStudy (anonymous):

6(1-(3/4) = 24

OpenStudy (anonymous):

?

hartnn (hartnn):

u forgot 1 /

hartnn (hartnn):

6/(1-(3/4) = 24

OpenStudy (anonymous):

Bah, I knew there had to be a fraction in there . . .

hartnn (hartnn):

6/(1-(3/4)) = 24

OpenStudy (anonymous):

oh yeah.. missed the first divisor.. thanks hartnn

OpenStudy (andriod09):

hart is right. wow.

OpenStudy (anonymous):

Didn't think to subtract from one yet, though. Good one!

OpenStudy (anonymous):

learnd that in my ics class for operators in java haha

hartnn (hartnn):

that was good one @Nameless

OpenStudy (anonymous):

thanks, and thanks for the help if u helped me :) see yall soon